# Exact inference via quasi-conjugacy in two-parameter Poisson–Dirichlet hidden Markov models

### v1

@@ 3.2 Posterior inference: filtering, smoothing, and forecasting
Direct use of the transition dynamics (9) is intractable. We instead exploit a dual representation of the PD diffusion in terms of a pure-death process $N(t)$ with finite starting point (griffiths2024dual). This duality allows the transition $X(t_{k})\mid X(t_{k-1})$ to be expressed, conditionally on observed data, through a latent partition with finite support, yielding tractable recursive filtering.
We now summarize the main inferential results for the two-parameter Poisson–Dirichlet HMM (8), with full details in the Appendix (Sections A.4.1-A.4.2).
@@
All posterior laws of interest—including filtered, smoothed, and interpolated distributions—belong to the following finite mixture family:
$\mathcal{F}:=\left\{\nu=\sum_{\lambda\in\Lambda}w_{\lambda}\mathrm{PD}_{\alpha,\theta}^{\lambda}:\ \Lambda\subset\mathcal{P},\ |\Lambda|<\infty,\ \sum w_{\lambda}=1\right\},$ (11)
@@
where $w_{\lambda}$ has finite support and reflects the full observation history. The evolution of these latent quantities (sets $\Lambda_{t}$ and weights $w_{\lambda}(t)$) is governed explicitly by the dual process and the coagulation algebra introduced in Section 2.3.
@@
i.e., the set of partitions that can be obtained by removing objects from partitions in $\Lambda$.
@@
    where $\hat{v}_{\omega}\propto v_{\omega}\mathrm{CRP}_{\omega\uparrow\pi^{k}}$, $\mathrm{CRP}_{\omega\uparrow\pi^{k}}$ is as in (7), and $\mathrm{Coag}_{\alpha,\theta}(\omega,\pi^{k})$ is fully described in (34).
@@
Fig. 4 illustrates this process: an initial observed parition $\pi^{0}=(2,1)$ (left graph) yields posterior $X(t_{0})\sim\mathrm{PD}_{\alpha,\theta}^{(2,1)}$. The prediction step computes a lower set $L((2,1))$ via the dual death process (center graph), forming the support of the latent partitions in $\nu_{1|0}$. Then, observing $\pi^{1}=(1,1)$ leads to the updated support $\mathrm{coag}(L((2,1)),(1,1))$ (right graph) through the product expansion. The recursion continues analogously at later times.
This structure enables fast and tractable online inference: filtering requires only propagation and update of a finite set of partitions $\lambda$ and their weights $w_{\lambda}$, with no need for simulation or numerical integration.
@@
The key insight is that smoothing retains the same latent partition structure as filtering: at any time $t_{k}$, the smoothed distribution of $X(t_{k})$ remains a finite mixture over conditionally conjugate components $\mathrm{PD}_{\alpha,\theta}^{\mu}$, where each index $\mu$ reflects a plausible state of the latent partition consistent with both past and future data. As in filtering, the latent partition act as sufficient statistics for posterior inference, and their evolution is governed by algebraic operations on partitions.
@@
Despite the bidirectional nature of the algorithm, tractability is fully preserved. All steps involve only finite sets of partitions, and weights can be computed exactly using closed-form expressions. No simulation or numerical integration is required. Full mathematical details, including the mixture representation and formal proof of correctness, are provided in Theorem 4 of the Appendix.
@@
The same recursive structure extends naturally to inference at unobserved time points. For forecasting beyond the last observation time $t_{N}$, the distribution of the latent state $X(t)$ for $t>t_{N}$ is obtained by propagating the filtered distribution $\nu_{N|0:N}$. Thanks to Theorem 2 in the Appendix, this again takes the form (12), where the weights $w_{\lambda}(t)$ reflect the evolution of the dual process beyond the observation window. E.g., if at $t_{N}$ the latent partition has support as in the rightmost graph of Fig. 4, call it $\Lambda_{0:N}$, then $\Lambda_{t}$ in (12) is the lower set $L(\Lambda_{0:N})$, which in this case is the union of $\Lambda_{0:N}$ with $\{(3),(2),(1),\emptyset\}$. As the forecast horizon grows ($t\to\infty$), the probability mass over $L(\Lambda_{0:N})$ progressively concentrates towards node in a neighborhood of $\emptyset$, implying that influence of past data vanishes and the forecast distribution converges to the stationary law $\mathrm{PD}_{\alpha,\theta}$ (cf. Corollary 1)
@@
$\Pi^{\prime}(t)|\lambda(t)=\lambda\sim\mathrm{CRP}_{\lambda\uparrow\bullet},\quad\lambda(t)|\Pi^{0:N}\sim w_{\bullet}(t)$ (18)
with $\mathrm{CRP}_{\lambda\uparrow\bullet}$ as in (7). In particular, if $t>t_{N}$, this is achieved by evolving the current sufficient statistic $\lambda(t_{N})$ forward in time through the dual process to obtain $\{w_{\lambda}(t),\lambda\in\Lambda_{t}\}$. Then, conditional on $\lambda(t)=\lambda$, a future observation $\Pi^{\prime}(t)$ has law $\mathrm{CRP}_{\lambda\uparrow\bullet}$. This yields interpretable predictions and simulations of future observations.
More generally, our framework enables exact posterior predictive inference for functionals of the latent measure $X(t)$, such as entropy or heterozygosity, i.e., the probability that two randomly chosen individuals belong to different types. This is obtained by applying explicit moment formulas for $\mathrm{PD}_{\alpha,\theta}$ laws. For instance, heterozygosity corresponds to the second moment of $X(t)$, which admits a closed-form expression. This allows principled forecasting and uncertainty quantification for summary statistics of interest, without requiring forward simulation or numerical integration.
@@
All mixture components and weights in Theorem 1 are available in closed form and detailed in the Appendix, as well as algorithms for exact simulation and evaluation (cf. Algorithms 38).
Theorem 1 establishes that posterior inference for the hidden model (8) is tractable: all relevant conditional distributions admit closed-form expressions as finite mixtures of quasi-conjugate components, with no need for Markov chain Monte Carlo or Sequential Monte Carlo approximations. This result enables exact recursive inference for infinite-dimensional nonparametric latent dynamics, using only combinatorial structures on partitions. Fig. 6 illustrates the resulting inferential procedures, which unify prediction, filtering, smoothing, and interpolation within a single coherent framework.
@@ 4.1 Summary of the algorithms
The full filtering and smoothing procedures, including intermediate computational steps, are detailed in the pseudocodes provided in Section A.5. In particular, Algorithms 1 and 2 serve as auxiliary routines for simulating from an $\varepsilon$-truncated $\mathrm{PD}_{\alpha,\theta}$ distribution and a conditional $\varepsilon$-truncated $\mathrm{PD}_{\alpha,\theta}^{\pi}$, respectively. The core inferential operations are handled by the remaining algorithms: Algorithm 3 samples from the propagation step (cf. (16)); Algorithm 5 covers the update step (cf. (44)); and Algorithm 6 combines these components to perform online filtering via the recursive operations (cf. (41)). For offline inference, Algorithms 7 and 8 support latent state estimation and partition simulation, respectively, after observing the full dataset.
@@ 4.2 Computability and algorithmic strategies
When initialized at $X(0)\sim\mathrm{PD}_{\alpha,\theta}$, the latent system satisfies computability in the sense of chaleyat2006computable: all conditional distributions of $X(t)$ lie in the finite mixture family $\mathcal{F}$ of Equation (11), and thus admit exact evaluation without resorting to stochastic truncation, MCMC, or SMC.
Nevertheless, efficient approximations are often desirable. Two key aspects are addressed in our implementation. First, computing the mixture weights $p^{\downarrow}_{\lambda,\bullet}(t)$ in Equation (42)—the transition probabilities of a pure-death process on $\mathcal{P}$—can be computationally intensive. Although exact formulas are available, we use a Monte Carlo approach based on the Gillespie algorithm (gillespie2007stochastic) to simulate paths of the death process (see Algorithm 4). This approximation may encounter numerical difficulties when starting from large partitions and propagating over short time intervals; in such cases, Gaussian approximations (griffiths1984asymptotic) provide a useful alternative.
@@
where $\mathcal{S}_{\kappa}(l,u,y)=(u-l)+\frac{2}{\kappa}(l-y)\mathbbm{1}\{y<l\}+\frac{2}{\kappa}(y-u)\mathbbm{1}\{y>u\}$, and $l,u$ are the $\kappa/2$ and $1-\kappa/2$ quantiles of the filtering distribution for a functional quantity $y$. This score measures forecast accuracy by penalizing both the width of the predictive interval and any observations that fall outside it, with lower values indicating better performance. We apply this to the heterozygosity $H_{2}(x):=1-\sum_{j\geq 1}x_{j}^{2}$, comparing its true value with the 95% credible intervals from the filter. Our results support pruning as a valid and efficient approach for practical implementation.
@@ 4.3 Parameter estimation
The marginal likelihood of the observed data sequence is given by
$\Pr(\Pi^{0}=\pi^{0},\dots,\Pi^{N}=\pi^{N})=\mathrm{PSF}_{\alpha,\theta}(\pi^{0})\prod_{k=1}^{N}\sum_{\omega\in L(\Lambda_{0:k-1})}v_{\omega}\mathrm{CRP}_{\omega\uparrow\pi^{k}}$ (20)
@@ 4.4 Numerical experiments
We illustrate our methodology using synthetic data from a known $\mathrm{PD}_{\alpha,\theta}$ diffusion trajectory. The goal is to estimate the heterozygosity functional of the latent state $x$, given by $H_{2}(x):=1-\sum_{j\geq 1}x_{j}^{2}$. Although we focus on $H_{2}$, our approach can be readily adapted to other functionals on the ordered infinite simplex, such as entropy and concentration indices (patil1982diversity).
@@
Model parameters are estimated by maximizing the marginal likelihood over a discrete grid, yielding $(\hat{\alpha},\hat{\theta})=(0,1.5)$—slightly underestimating $\alpha$. Fig. 8 displays marginal posterior estimates for $H_{2}(x)$ at the observed time points. Violin plots represent posterior densities (via kernel density estimation), triangles mark 95% credible intervals, and white diamonds indicate the observed heterozygosities $\hat{H}_{2}(\pi^{k})=1-\sum_{j}(\pi_{j}^{k}/|\pi^{k}|)^{2}$. Both the filtering and smoothing procedures (top panels) successfully track the true heterozygosity trajectory (black line), even under this slight model mismatch. See Section A.5.3 for further plots.
Notably, our duality-based filter achieves superior accuracy compared to the Bootstrap Particle Filter (BPF)—a leading benchmark in the literature (chopin2020introduction). While the BPF (bottom left panel) performs reasonably well, it requires expensive simulation of $\mathrm{PD}_{\alpha,\theta}$-diffusion paths (Eq. (9)) and exact importance weight computations (Eq. (4)) for each particle and time step, leading to substantial computational overhead. Furthermore, as seen in Fig. 8, the BPF may occasionally yield posteriors with greater variance and pronounced multimodality, especially at intermediate time points (cf. also Section 4.5).
@@
Finally, as shown in the bottom right panel, inference based on independent priors—that is, ignoring temporal dependence—may incidentally approximate the true heterozygosity but lacks robustness and statistical adequacy in dynamic or data-sparse regimes. It also precludes interpolation at arbitrary time points, which our model handles naturally and with no additional effort.
@@ 4.5 Application to dynamic face-to-face interaction networks
Each day’s data yield a time-indexed network in which nodes represent individuals and edges denote interactions. We focus on a single day—June 28, 2009—aggregating the data into 30-minute windows to produce 12 networks. The connected components of each graph induce an unlabelled partition representing clusters of individuals in contact. Fig. 1 in Section 1 displays four representative networks, each inducing partitions $(11,9,4,3)$, $(38,2)$, $(22,3,2,2,2)$, and $(17,4,2,2,2)$.
We model the time-varying group structure using a two-parameter Poisson–Dirichlet diffusion for the latent probabilities of cluster sizes. Our inferential target is the heterozygosity functional $H_{2}(x):=1-\sum_{j\geq 1}x_{j}^{2}$, which approximates the probability that two randomly chosen individuals belong to different components, serving as a proxy for effective population mixing and epidemic spread.

### v2

@@ 3.2 Posterior inference: filtering, smoothing, and forecasting
Direct use of the transition dynamics (10) is intractable. We instead exploit a dual representation of the two-parameter Poisson–Dirichlet diffusion in terms of a pure-death process $N(t)$ with finite starting point (griffiths2024dual). This duality allows the transition $X(t_{k})|X(t_{k-1})$ to be expressed, conditionally on observed data, through a latent partition with finite support, yielding tractable recursive filtering.
We now summarize the main inferential results for the two-parameter Poisson–Dirichlet HMM (9), with full details in the Supplement (Sections A.3.1-A.3.2).
@@
All posterior laws of interest, including filtered, smoothed, and interpolated distributions, belong to the following finite mixture family:
$\mathcal{F}:=\left\{\nu=\sum_{\lambda\in\Lambda}w_{\lambda}\mathrm{PD}_{\alpha,\theta}^{\lambda}:\ \Lambda\subset\mathcal{P},\ |\Lambda|<\infty,\ \sum_{\lambda\in\Lambda}w_{\lambda}=1\right\},$ (12)
@@
where $w_{\lambda}$ has finite support and reflects the full observation history. The evolution of these latent quantities (sets $\Lambda_{t}$ and weights $w_{\lambda}(t)$) is governed explicitly by the dual process and the coagulation algebra introduced in Section $2.$
@@
i.e., the set of partitions that can be obtained by removing units from partitions in $\Lambda$.
@@
    where $\hat{v}_{\omega}\propto v_{\omega}\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})$, $\mathrm{PSF}_{\alpha,\theta}^{\omega}$ is as in (4), and $\mathrm{Coag}_{\alpha,\theta}(\omega,\pi^{k})$ is fully described in (37).
@@
Fig. 4 illustrates this process: an initial observed partition $\pi^{0}=(2,1)$ (left graph) yields posterior $X(t_{0})\sim\mathrm{PD}_{\alpha,\theta}^{(2,1)}$. The prediction step computes a lower set $L((2,1))$ via the dual death process (center graph), forming the support of the latent partitions in $\nu_{1|0}$. Then, observing $\pi^{1}=(1,1)$ leads to the updated support $\mathrm{coag}(L((2,1)),(1,1))$ (right graph) through the product expansion. The recursion continues analogously at later times.
This structure enables tractable online inference: filtering requires only propagation and update of a finite set of partitions $\lambda$ and their weights $w_{\lambda}$, with no need for simulation or numerical integration.
@@
The key insight is that smoothing retains the same latent partition structure as filtering: at any time $t_{k}$, the smoothed distribution of $X(t_{k})$ remains a finite mixture over conditionally conjugate components $\mathrm{PD}_{\alpha,\theta}^{\mu}$, where each index $\mu$ reflects a plausible state of the latent partition consistent with both past and future data. As in filtering, the latent partition acts as sufficient statistics for posterior inference, and their evolution is governed by algebraic operations on partitions.
@@
Despite the bidirectional nature of the algorithm, tractability is fully preserved. All steps involve only finite sets of partitions, and weights can be computed exactly using closed-form expressions. No simulation or numerical integration is required. Full mathematical details, including the mixture representation and formal proof of correctness, are provided in Theorem 4 of the Supplement.
@@
The same recursive structure extends naturally to inference at unobserved time points. For forecasting beyond the last observation time $t_{N}$, the distribution of the latent state $X(t)$ for $t>t_{N}$ is obtained by propagating the filtered distribution $\nu_{N|0:N}$. Thanks to Theorem 2 in the SM, this again takes the form (13), where the weights $w_{\lambda}(t)$ reflect the evolution of the dual process beyond the observation window. E.g., if at $t_{N}$ the latent partition has support as in the rightmost graph of Fig. 4, call it $\Lambda_{0:N}$, then $\Lambda_{t}$ in (13) is the lower set $L(\Lambda_{0:N})$, which in this case is the union of $\Lambda_{0:N}$ with $\{(3),(2),(1),\emptyset\}$. As the forecast horizon grows ($t\to\infty$), the probability mass over $L(\Lambda_{0:N})$ progressively concentrates towards nodes in a neighborhood of $\emptyset$, implying that the influence of past data vanishes and the forecast distribution converges to the stationary law $\mathrm{PD}_{\alpha,\theta}$ (cf. Corollary 1)
@@
$\Pi^{\prime}(t)|\lambda(t)=\lambda\sim\mathrm{PSF}_{\alpha,\theta}^{\lambda},\quad\lambda(t)|\Pi^{0:N}\sim w_{\bullet}(t)$ (19)
with $\mathrm{PSF}_{\alpha,\theta}^{\lambda}$ as in (4). In particular, if $t>t_{N}$, this is achieved by evolving the current sufficient statistic $\lambda(t_{N})$ forward in time through the dual process to obtain $\{w_{\lambda}(t),\lambda\in\Lambda_{t}\}$. Then, conditional on $\lambda(t)=\lambda$, a future observation $\Pi^{\prime}(t)$ has law $\mathrm{PSF}_{\alpha,\theta}^{\lambda}$. This yields interpretable predictions and simulations of future observations.
The finite-mixture representation (13) also supports principled forecasting and uncertainty quantification for functionals of the latent random measure $X(t)$, such as diversity indices (patil1982diversity). For example, heterozygosity has a natural interpretation as the probability that two samples drawn from $X(t)$ belong to different types. Under the finite-mixture representation, the posterior distribution of heterozygosity at time $t$ is itself a finite mixture of the corresponding distributions induced by the $\mathrm{PD}_{\alpha,\theta}^{\lambda}$ laws, from which sampling is straightforward. This strategy is implemented in Section 4.4.
@@
Theorem 1 establishes that posterior inference for the hidden Markov model (9) is tractable: all relevant conditional distributions admit closed-form finite-mixture representations of quasi-conjugate components. The associated mixture weights and simulation algorithms are detailed in the Supplement (cf. Algorithms 38). This yields exact recursive inference for infinite-dimensional nonparametric latent dynamics without resorting to MCMC or SMC methods. Fig. 6 illustrates the resulting framework, which unifies prediction, filtering, smoothing, and interpolation.
@@ 4.1 Summary of the algorithms
The full filtering and smoothing procedures, together with all intermediate computational steps, are detailed in the pseudocode provided in Section A.4. In particular, Algorithms 1 and 2 are auxiliary routines for simulating from an $\varepsilon$-truncated $\mathrm{PD}_{\alpha,\theta}$ distribution and from a conditional $\varepsilon$-truncated $\mathrm{PD}_{\alpha,\theta}^{\pi}$, respectively. Here, $\varepsilon$-truncation refers to terminating the stick-breaking construction once the remaining (residual) mass falls below $\varepsilon$, thereby yielding a finite-dimensional, ordered approximation of the Poisson–Dirichlet law. The core inferential operations are handled by the remaining algorithms: Algorithm 3 samples from the propagation step (cf. (17)); Algorithm 5 covers the update step (cf. (47)); and Algorithm 6 combines these components to perform online filtering via the recursive operations (cf. (44)). For offline inference, Algorithms 7 and 8 support latent state estimation and partition simulation, respectively, after observing the full dataset.
@@ 4.2 Computability and algorithmic strategies
When initialized at $X(0)\sim\mathrm{PD}_{\alpha,\theta}$, the latent system satisfies computability in the sense of chaleyat2006computable: all conditional distributions of $X(t)$ lie in the finite mixture family $\mathcal{F}$ of Equation (12), and thus admit exact evaluation without resorting to MCMC or SMC.
Nevertheless, efficient approximations are often desirable. Two key aspects are addressed in our implementation. First, computing the mixture weights $p^{\downarrow}_{\lambda,\bullet}(t)$ in Equation (45) (the transition probabilities of a pure-death process on $\mathcal{P}$ related to Kingman’s coalescent in Eq. (10)) can be computationally intensive. Although exact formulas are available, we use a Monte Carlo approach based on the Gillespie algorithm (gillespie2007stochastic) to simulate paths of the death process (see Algorithm 4). This approximation may encounter numerical difficulties when starting from large partitions and propagating over short time intervals; in such cases, Gaussian approximations (griffiths1984asymptotic) provide a useful alternative.
@@
where $\mathcal{S}_{\kappa}(l,u,y)=(u-l)+\frac{2}{\kappa}(l-y)\mathbbm{1}\{y<l\}+\frac{2}{\kappa}(y-u)\mathbbm{1}\{y>u\}$, and $l,u$ are the $\kappa/2$ and $1-\kappa/2$ quantiles of the filtering distribution for a functional quantity $y$. This score measures forecast accuracy by penalizing both the width of the predictive interval and any observations that fall outside it, with lower values indicating better performance. We apply this to the heterozygosity $H_{2}(x):=1-\sum_{i\geq 1}x_{i}^{2}$, comparing its true value with the 95% credible intervals from the filter. Our results support pruning as a valid and efficient approach for practical implementation.
@@ 4.3 Parameter estimation
The marginal likelihood (5) of the observed data sequence is given by
$\Pr(\Pi^{0}=\pi^{0},\dots,\Pi^{N}=\pi^{N})=\mathrm{PSF}_{\alpha,\theta}(\pi^{0})\prod_{k=1}^{N}\sum_{\omega\in L(\Lambda_{0:k-1})}v_{\omega}\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})$ (21)
@@ 4.4 Numerical experiments
We illustrate our methodology using synthetic data from a known $\mathrm{PD}_{\alpha,\theta}$ diffusion trajectory. The goal is to estimate the heterozygosity functional of the latent state $x$, given by $H_{2}(x):=1-\sum_{i\geq 1}x_{i}^{2}$. Although we focus on $H_{2}$, our approach can be readily adapted to other functionals on the ordered infinite simplex, such as entropy and diversity indices (patil1982diversity).
@@
Model parameters are estimated by maximizing the marginal likelihood over a discrete grid, yielding $(\hat{\alpha},\hat{\theta})=(0,1.5)$—slightly underestimating $\alpha$. Fig. 8 displays marginal posterior estimates for $H_{2}(x)$ at the observed time points. Violin plots represent posterior densities (via kernel density estimation), triangles mark 95% credible intervals, and white diamonds indicate the observed heterozygosities $\hat{H}_{2}(\pi^{k})=1-\sum_{i}(\pi_{i}^{k}/|\pi^{k}|)^{2}$. Both the filtering and smoothing procedures (top panels) successfully track the true heterozygosity trajectory (black line), even under this slight model mismatch. See Section A.4.3 for further plots.
Notably, our duality-based filter achieves superior accuracy compared to the Bootstrap Particle Filter (BPF), a leading benchmark in the literature (chopin2020introduction). While the BPF (bottom left panel) performs reasonably well, it requires expensive simulation of $\mathrm{PD}_{\alpha,\theta}$-diffusion paths (Eq. (10)) and exact importance weight computations (Eq. (6)) for each particle and time step, leading to substantial computational overhead. Furthermore, as seen in Fig. 8, the BPF may occasionally yield posteriors with greater variance and pronounced multimodality, especially at intermediate time points (cf. also Section 4.5).
@@
Finally, as shown in the bottom right panel, inference based on independent priors (ignoring temporal dependence) may incidentally approximate the true heterozygosity but lacks robustness and statistical adequacy in dynamic or data-sparse regimes. It also precludes interpolation at arbitrary time points, which our model handles naturally and with no additional effort.
@@ 4.5 Application to dynamic face-to-face interaction networks
Each day’s data yield a time-indexed network in which nodes represent individuals and edges denote interactions. We focus on a single day, June 28, 2009, aggregating the data into 30-minute windows to produce 12 networks. The connected components of each graph induce an unlabelled partition representing clusters of individuals in contact. Fig. 1 in Section 1 displays four representative networks, each inducing partitions $(11,9,4,3)$, $(38,2)$, $(22,3,2,2,2)$, and $(17,4,2,2,2)$.
We model the time-varying group structure using a two-parameter Poisson–Dirichlet diffusion for the latent probabilities of cluster sizes. Our inferential target is the heterozygosity functional $H_{2}(x):=1-\sum_{i\geq 1}x_{i}^{2}$, which approximates the probability that two randomly chosen individuals belong to different components, serving as a proxy for effective population mixing and epidemic spread.